1. A 0.2µF capacitor has a charge of 20 µC. Find the voltage and energy.


2. If the energy stored in a 1/8 F capacitor is 25 J, find the voltage and charge.


3. Find the maximum and minimum values of capacitance that can be obtained 

from ten 1µF capacitors.


4. Suppose you have a 9V battery, a 2μF capacitor, and a 7.40μF capacitor. 

(a) Find the charge and energy stored if the capacitors are connected to the 

battery in series. (b) Do the same for a parallel connection.


5. In an open-heart surgery, a much smaller amount of energy will defibrillate 

the heart. (a) What voltage is applied to the 8µF capacitor of a heart 

defibrillator that stores 40 J of energy? (b) Find the amount of stored charge.


6. If the current through a 1mH inductor is 𝑖(௧) = 20 cos 100𝑡 𝑚𝐴, find the 

terminal voltage and the energy stored.


7. Find the maximum and minimum values of inductance that can be obtained 

using ten 10mH inductors.


1
Expert's answer
2021-12-15T04:01:51-0500

1.



"Q=CU=>U=\dfrac{Q}{C}=\dfrac{20\times10^{-6}C}{0.2\times10^{-6}F}=100V"





"E=\dfrac{Q^2}{2C}=\dfrac{(20\times10^{-6}C)^2}{2(0.2\times10^{-6}F)}=0.001J"

2.



"E=\dfrac{Q^2}{2C}=>Q=\pm\sqrt{2CE}""Q=\pm\sqrt{2(\dfrac{1}{8}F)(25J)}=\pm2.5C""U=\dfrac{Q}{C}=\dfrac{\pm 2.5C}{\dfrac{1}{8}F}=\pm20V"

3.

Series capacitors



"\dfrac{1}{C_{TS}}=\dfrac{1}{C_1}+\dfrac{1}{C_2}+...+\dfrac{1}{C_{10}}""=\dfrac{1}{10C_1}=\dfrac{1}{10\times 10^{-6}F}""C_{TS}=10^5F"

Parallel capacitors



"C_{TP}=C_1+C_2+...+C_{10}=10C_1""=10\times 10^{-6}F=10^{-5}F"


The maximum value of capacitance that can be obtained is "10^5F."

The minimum value of capacitance that can be obtained is "10^{-6}F" (when we use only one capacitor).

If we have to use all ten 1µF capacitors together then the maximum value of capacitance that can be obtained is "10^5F," and the minimum value of capacitance that can be obtained is "10^{-5}F."


4.

(a) Series capacitors



"\dfrac{1}{C_{TS}}=\dfrac{1}{C_1}+\dfrac{1}{C_2}=\dfrac{1}{2\times 10^{-6}F}+\dfrac{1}{7.4\times 10^{-6}F}""=\dfrac{4.7}{7.4\times 10^{-6}F}""Q_S=C_{TS}U=\dfrac{7.4\times 10^{-6}F}{4.7}\cdot9V""\approx14.17\times 10^{-6} C=14.17\mu C""E_S=\dfrac{C_{TS}U^2}{2}=\dfrac{7.4\times 10^{-6}\cdot(9V)^2}{2(4.7)}""\approx63.766\times 10^{-6} J=63.766\mu J"

(b) Parallel capacitors



"C_{TP}=C_1+C_2=2\times 10^{-6}F+7.4\times 10^{-6}F""=9.4\times 10^{-6}F""Q_P=C_{TP}U=9.4\times 10^{-6}F\cdot9V""=84.6\times 10^{-6} C=84.6\mu C""E_P=\dfrac{C_{TP}U^2}{2}=\dfrac{9.4\times 10^{-6}\cdot(9V)^2}{2}""=0.3807\times 10^{-3} J=0.3807m J"



5.

(a)



"E=\dfrac{CU^2}{2}=>U=\pm\sqrt{\dfrac{2E}{C}}""=\pm\sqrt{\dfrac{2(40J)}{8\times 10^{-6}F}}=\pm(\sqrt{10}\times10^3)V"

(b)


"Q=CU=\pm(\sqrt{10}\times10^3)V\cdot(8\times 10^{-6}F)""=\pm(8\sqrt{10}\times10^{-3} )C"

6.

i.



"V(t)=L\dfrac{di}{dt}=10^{-3}(-20(100)\times10^{-3}\sin(100t))V""=-2\sin(100t)mV"

ii.



"E=\dfrac{Li^2}{2}=\dfrac{10^{-3}(20\cos(100t)\times10^{-3})^2}{2}J""=0.2\cos^2(100t)\mu J"


7.

Series inductances



"L_{TC}=L_1+L_2+...+L_{10}""=10(0.01H)=0.1H"

Parallel inductances



"\dfrac{1}{L_{TP}}=\dfrac{1}{L_1}+\dfrac{1}{L_2}+...+\dfrac{1}{L_{10}}""=10(\dfrac{1}{0.01H})""L_{TP}=\dfrac{0.01H}{10}=0.001H=1mH"

The maximum possible inductance is "0.1H=100mH."

The minimum possible inductance is "0.001H=1mH."